There are 3 consecutive odd integers. Three times the largest is seven times the smallest. What are the integers?
Answer by oberobic(2304)

3 consecutive odd integers can be called: x, x+2, and x+4. We are told 3(x+4) = 7x, which is simply 3x + 12 = 7x. Subtracting 3x from both sides, you have 12 = 4x. Dividing through by 4, you have x = 3.
ALWAYS read the question carefully. It asks what are the integers, so the answer is: 3, 5, and 7.
ALWAYS check. Is 3*7= 7*3? Yes. Check.